OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (2, -2) with 0° ≤ θ < 360

OpenStudy (tkhunny):

Jut pick them out! Polar coordinates look like this: \((r,\theta)\) I'll do the hard one. You do the easy one. \((-2\sqrt{2},3\pi/4)\;or\;(-2\sqrt{2},135º)\)

OpenStudy (anonymous):

wait im sorry what

OpenStudy (anonymous):

@hartnn you have any idea?

hartnn (hartnn):

x=2,y=-2 r = \sqrt (x^2+y^2) theta = arctan(y/x)

OpenStudy (anonymous):

ok what does that mean xD?

hartnn (hartnn):

i just gave you formulas, you had a point x,y = 2,-2 to convert it into polar , we use this formulas r = \sqrt (x^2+y^2) theta = arctan(y/x) and then polar co-ordinate will be r, theta